if M,N and T are in arithmetic progression prove that [M+2N-T][2N+T-M][T+M-N]=4MNT

Grade:10

1 Answers

Santhosh
6 years ago
Put M=a, N=a+d T=a+2d and then by solving given lhs we get (2a)(2a+4d)(a+d) = 4a(a+d)(a+2d) =4MNT Provided the assumption is because given they are in AP If they are in GP use M=a÷r ,N=a,T=ar for else question